OpenStudy (anonymous):

What equation is perpendicular to y= -x/2-4?

OpenStudy (anonymous):

y= 2x-4

OpenStudy (anonymous):

use the theory that the product of slope od 2 perpendicluar lines is -1

OpenStudy (anonymous):

slope of the given equation is -1/2 lines perpendicular to each other => multiply their slopes we get -1 therefore the slope of the required line is 2 so, the equation is y=2x-4

OpenStudy (anonymous):

y=2x-4 is not the correct answer.

OpenStudy (anonymous):

the slope of the perpendicular line is always negative reciprocal of the slope of original line. your given slope is -1/2 it's reciprocal is -2/1 or simply -2 and multiply by negative so it become 2. then substitute to the equation. y = 2x - 4

OpenStudy (anonymous):

y=2x+4

OpenStudy (anonymous):

COCOTSOI IS RIGHT WITH THIS ONE

OpenStudy (anonymous):

no there is a smal problem in the question itself u have mentioned the line will be perpendicular but not what will be its intercept.Which means that the eqn of any line perpendicular to it is Y=2X+c'(where the value of C' will be determined by any coordinate on the new line )

OpenStudy (anonymous):

the correct answer is either y=-4x-1/2, y=-1/2x+4, y=2x+2, or y=1/2x+4 ? Just not sure which of these is correct .

OpenStudy (anonymous):

y=2x+2 as it is the only line with slope 2 There are more than one answer if there are no choices given

OpenStudy (anonymous):

the + or - doesn't matter at all. it will just move the line horizontally, the meaning of perpendicular is it makes a 90 degree angle when it intersect to the other line that's what we call perpendicular line. ok??

OpenStudy (anonymous):

 all are perpendicular

OpenStudy (anonymous):

What is the slope of a vertical line?

OpenStudy (anonymous):

undefined,

OpenStudy (anonymous):

1/0 which is undefined

OpenStudy (anonymous):

or in some books it is infinite.

OpenStudy (anonymous):

KUMAR SAJAN COCO -SIGN brings A VERY BIG CHANGE IN MATHS so its matters everywhere u cant say it doesnt matters and i told u guys cocotsoi method is right and seargent pepper the answer cocotsoi gave above is the right one or else ur question is wrong

OpenStudy (anonymous):

kumar sajn 1/0 answer is infinity

OpenStudy (anonymous):

The slopes of parallel lines sum to 0, are equal, have negative reciprocals, or sum to 1?

OpenStudy (anonymous):

slope of //lines are equal

OpenStudy (anonymous):

the only question is what is the perpendicular equation. it is: y = 2x + b -----> where b can be any integer. it can be: y = 2x + 1 y = 2x + 2 y = 2x + 3 y = 2x + 4 y = 2x + 5 or y = 2x - 1 y = 2x - 2 y = 2x - 3 y = 2x - 4 y = 2x -5

OpenStudy (anonymous):

what is the slope of a line that contains the points, (4,6) & (12,3) ?

OpenStudy (anonymous):

(3-6)/(12-4) -3/8

OpenStudy (anonymous):

use the fromula m=y2-y1/x2-x1

OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y=4x+5? y=-1/4x y=-4x+5 y=4x-3 y=5x+4

OpenStudy (anonymous):

but that's not one of the answers given, so would it possibly be y=-4x+5?

OpenStudy (anonymous):

y=4x-3

OpenStudy (anonymous):

@WAHEEDAHMED do you know what are you saying? please don't teach if you are not sure about it.

OpenStudy (anonymous):

@kumarsajan thank you so much for your help, could you help me with another problem like this real quick? (:

OpenStudy (anonymous):

in parallel lines the slope is always equal. remember if 2 lines are parallel they will never intersect ( never meet at any point in the graph). as i said before y = mx + b ---> change in b will only move your line in the graph. if it is still the same as WAHEEDAHMED said then you are drawing a line above the original line which are not parallel.

OpenStudy (anonymous):

Which of the following is the equation of a line that is parallel to the line 3x-2y=4? y=-2/3x-4 y=2/3x+4 y=3/2x y=-3/2x-2

OpenStudy (anonymous):

hey kumar sajan i didnt saw the answer y=4x-3 or i would have told him the same stop acting as the smartest guy in this world i did it by mistake and every post i was also proving him that the slope for the parallel will be the same that is 4 so back off a littlebit smarty pants

OpenStudy (anonymous):

@WAHEEDAHMED I"M NOT A GUY !!!!

OpenStudy (anonymous):

@WAHEEDAHMED dude read this, KUMAR SAJAN COCO -SIGN brings A VERY BIG CHANGE IN MATHS so its matters everywhere u cant say it doesnt matters and i told u guys cocotsoi method is right and seargent pepper the answer cocotsoi gave above is the right one or else ur question is wrong so can you explain me what is wrong in my answer above. :))

OpenStudy (anonymous):

i was not to u sergeant

OpenStudy (anonymous):

one minute let me have a look at it ....

OpenStudy (anonymous):

Can you guys please help me with the problem I posted above @WAHEEDAHMED comment?

OpenStudy (anonymous):

first you will change the equation to slope-intercept form ---> y = mx + b 3x-2y=4 --> devide all by -2 (3x - 2y = 4)/-2 -3/2x + y = -2 ----> transpose -3/2x to the other side. y = 3/2x -2 so what do you think can be the answer. :))

OpenStudy (anonymous):

remember the slope of parallel lines are always equal, y = mx + b ---> where m is the slope. and b is any integer.

OpenStudy (anonymous):

What is the standard form equation of the line y=x/5-2?

OpenStudy (anonymous):

slope of the given equation is -1/2 lines perpendicular to each other => multiply their slopes we get -1 therefore the slope of the required line is 2 so, the equation is y=2x-4 so what was wrong in this answer isnt it right dude

OpenStudy (anonymous):

this answer was given by cocotsoi u can check it i was defending was it wrong

OpenStudy (anonymous):

Ohmygosh, y=2x-4 is not the correct answer for the millionth time, alright? Now get over it , my goodness....

OpenStudy (anonymous):

we are talking the method not the answer specificaly sergeant

OpenStudy (anonymous):

it is correct @sergeantpepper the only thing is that it is not in you multiple options. as i said before the b in y=mx + b doesn't matter if the line is perpendicular or not.

OpenStudy (anonymous):

so was i wrong supporting his answer kumar sajan now tell me

OpenStudy (anonymous):

i'm out of here. i think you can't find a good answer.

OpenStudy (anonymous):

no but you can't say that i'm wrong because my answer is also correct. :))@WAHEEDAHMED

OpenStudy (anonymous):

i dont wanna fight or am not saying your wrong or so

OpenStudy (anonymous):

i didnt say you are wrong anywhere if i have i am sorry about that